Please complete part 1, 2 and 3.

Given the following data 0 Find E (X) Find σ2 (3) Find σ x P (y 411 5-1-22 13 10 10 08 2 PX) 00
Answer #1

Answer

(A) E[x] is the expected value given as E[x] = \sum x*P(x) ​​​​​​

setting the data values, we get

=(4*0.17)+(5*0.22)+(6*0.19)+(7*0.13)+(8*0.11)+(9*0.10)+(10*0.08)

=0.68+1.10+1.14+0.91+0.88+0.90+0.80

=6.41

(B) Variance \sigma^2 is given as

\sigma^2 = \sum (x-\mu)^2*P(x)

setting the data values, we get

=((4-6.41)^2*0.17)+((5-6.41)^2*0.22)+((6-6.41)^2*0.19)+((7-6.41)^2*0.13)+((8-6.41)^2*0.11)+((9-6.41)^2*0.10)+((10-6.41)^2*0.08)

= 3.4819

(C) standard deviation \sigma is equal to square root of variance

\sigma = \sqrt{variance}

= \sqrt{3 . 4819}

=1.87